# Negate or Embrace: On How Misalignment Shapes Multimodal Representation Learning

### v1

@@ 3.2 Model Assumptions for Theoretical Analysis
We now present the technical assumptions underlying our theoretical analysis, based on the proposed LVM:
@@
Interpretation 3.1. Note that Eq. 5 essentially implies that, in each $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$, only a subset of semantic variables $A$ undergoes perturbations, regardless of the underlying causal structure among latent semantic variables. The rationale is that latent semantic variables can only be modified indirectly by altering observations, rather than through direct intervention, unless the latent causal structure is fully identified. In the text modality, it occurs through the misassignment of certain content words to specific image semantics during the captioning process. Unlike a direct intervention, this misalignment does not propagate to descendant semantic variables, as illustrated in Figure 2.
4 Theory: Identifiability Results of the Proposed Latent Variable Model
In this section, we theoretically analyze how misalignment impacts the identifiability of latent semantic variables in the proposed LVM, within the MMCL framework (§ 4.1). Based on these results, we further provide practical insights into how misalignment should be addressed in real-world applications (§ 4.2). Detailed proofs of the theoretical results are provided in App. C.
To begin, we restate the definition of block-identifiability [66] in the context of our problem setting:
@@ 4 Theory: Identifiability Results of the Proposed Latent Variable Model
4.1 Identifiability Result of Latent Semantic Variables under Misalignment
@@
Let $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$ be image-text pairs drawn from the data-generating process described in § 3, where $\mathbf{x}$ is generated according to Eq. 3 and $\mathbf{t}_{\theta}$ is generated by Eq. 4. Suppose that Asms. 3.1 and 3.2 hold. Denote by $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ the subset of semantic variables that preserved without bias in the text, and define its dimension as $n=|\mathbb{I}_{\theta}|-|\mathbb{I}_{\rho}|$. Let $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n}$ be sufficiently flexible, smooth functions. Then, Minimizing the loss $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ in Eq. 2 over samples $(\mathbf{x},\mathbf{t}_{\theta})$ guarantees that $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ block-identify the semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ in the sense of Defn. 4.1.
Discussion 4.1.
Thm. 4.1 formally establishes that, in the presence of misalignment, the unbiased semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ that are shared across modalities can be effectively recovered, up to a block-wise indeterminacy, by minimizing the MMCL objective. In contrast, components that are misaligned, specifically $\mathbf{s}_{\mathbb{I}_{\rho}}$ and $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$, are entirely excluded from the learned representations. We emphasize that this result holds regardless of any underlying latent graph structure among semantic variables. At its core, this result highlights the model’s capacity to focus exclusively on the aligned semantic aspects of the data. Furthermore, modality-specific, non-semantic factors, i.e., $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, are consistently discarded throughout the learning process. This further underscores the model’s ability to extract meaningful, cross-modal semantic information while filtering out semantically uninformative or noise-induced components.
@@ 4.2 Insights into Misalignment for Practice
Let the selection bias be $\theta=2^{n_{s}}-1$ and the perturbation bias be $\rho=1$, such that the full set of semantic variables $\mathbb{I}_{\mathbf{s}}$ is selected, and the perturbable semantic subset is trivial, i.e., $\mathbb{I}_{\rho}=\emptyset$. Then, all semantic variables $\mathbf{s}$ are block-identified via smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n_{s}}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n_{s}}$, when minimizing $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$.
Insight 4.1 (MMCL Pretraining on Large-Scale Data). Large-scale multimodal datasets (e.g., COCO [36], Conceptual Captions [59], LAION-5B [58]) often exhibit varying caption quality. Our analysis indicates that omitted or perturbed semantics are irretrievably lost in the learned representations, regardless of dataset size, although scale may mitigate sporadic misalignment by averaging its effects. Preserving a breadth of relevant semantic details is therefore crucial when pretraining foundation models, whose primary goal is to support diverse downstream tasks. As noted in Cor. 4.1, achieving this requires detailed and consistent annotation of image semantics. Consequently, improved caption control [34, 15, 13] is essential to avoid blind spots in semantic coverage.
@@
Consider an OOD setting in which a subset of semantic variables, $\mathbb{I}_{inv}\subset\mathbb{I}_{\mathbf{s}}$, remains invariant between training and testing environments, while the remaining semantic variables, $\mathbb{I}_{var}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{inv}$, undergo distribution shifts. If the union of omitted and perturbable semantic variables under selection bias $\theta$ and perturbation bias $\rho$ coincides with the environment-sensitive subset, i.e., $\mathbb{I}_{var}=\mathbb{I}_{\theta}^{c}\cup\mathbb{I}_{\rho},$ then the invariant semantic variables $\mathbf{s}_{\mathbb{I}_{inv}}$ are block-identified via smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{|\mathbb{I}_{inv}|}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{|\mathbb{I}_{inv}|},$ by minimizing $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$.
Insight 4.2 (Invariant Representation Learning). In tasks requiring robust OOD performance (e.g., domain generalization) [55, 2], semantic variables that are vulnerable to distribution shifts can undermine generalization. As noted in Cor. 4.2, misalignment may, counterintuitively, enhance robustness by selectively omitting or perturbing these vulnerable variables. This suggests that MMCL may offer a novel perspective on invariant representation learning [52, 30, 14], as auditing and curating text is more precise and interpretable, since language is distilled from human knowledge.
@@ 5.1 Numerical Simulation
We synthesize numerical data following the generative process described in § 3. Specifically, we sample modality-specific variables $\mathbf{m}_{x}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{x}})$ and $\mathbf{m}_{t}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{t}})$, each of dimension 5, along with full semantic variables $\mathbf{s}\sim\mathcal{N}(0,\Sigma_{\mathbf{s}})$, of dimension 10. Potential causal dependencies among these variables are encoded in their covariance matrices $\Sigma_{(\cdot)}$. To simulate various misalignment scenarios, we progressively increase the strength of selection and perturbation biases. For selection bias, we incrementally define selected subsets $\mathbb{I}_{\theta}$ as $\{1\},\dots,[10]$. [^5] For any positive integer $n>1$, we use the notation $[n]$ to denote the set $\{1,2,\dots,n\}$. For perturbation bias, we similarly use increasing subsets of $\mathbb{I}_{\rho}$, ranging from the empty set to $[9]$, applying additive Gaussian noise with a probability 0.75 for each semantic dimension $i\in\mathbb{I}_{\rho}$. These biases are applied separately: when analyzing selection bias, we set $\mathbb{I}_{\rho}=\emptyset$; conversely, when analyzing perturbation bias, we fix $\mathbb{I}_{\theta}=[9]$. Generation functions $\mathbf{g}_{x}$ and $\mathbf{g}_{t_{\theta}}$ are instantiated as randomly initialized invertible MLPs. Two modality-specific MLP encoders are trained for 100,000 steps using the $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ loss defined in Eq. 2, setting the representation dimension equal to the unbiased semantic dimensions. Further details on the experimental setup are provided in § D.1.
We conduct two main experiments. (i) Identification of semantics: We predict each dimension of the true semantic variables from the learned representations using a lightweight MLP, reporting the predictive $R^{2}$ score on holdout data. Results are averaged over three random seeds for each setting. (ii) Downstream performance: We evaluate the pretrained representations obtained under various bias conditions on downstream tasks. Specifically, we construct four regression task labels $y_{1},y_{2},y_{3}$, and $y_{4}$ by applying complex nonlinear functions to subsets of the semantic variables: $[3]$, $[5]$, $[7]$, and $[9]$, respectively. For classification, we binarize $y_{2}$ to produce binary labels. To evaluate the OOD generalization performance of the trained classifiers, we introduce a distribution shift in the observations $\mathbf{x}$ by applying a heavy-tailed transformation to dimensions $\{9,10\}$. Further details on the task design are provided in § D.1.
@@
The results in Figure 3 show that, under the independent latent variable scenario, unbiased semantic variables are clearly block-identified ($R^{2}\approx 1$), whereas misaligned semantics due to selection bias are effectively discarded ($R^{2}\approx 0$). In the dependent latent variable scenario, some misaligned semantics become partially predictable, reflecting inherent mutual predictability among strongly dependent variables [66, 76]. Modality-specific variables are consistently omitted from the representations across all settings. Similar effects are observed under perturbation bias, as illustrated in Figure 7. Notably, although our theoretical results hold only up to invertible mappings, simple linear regression already achieves high $R^{2}$ scores, as demonstrated in Figure 8. These findings consolidate the identifiability results in Thm. 4.1. Additional analyses on misassigned encoding dimensions and combined bias effects, are provided in § D.2.
@@
Downstream Performance.
As shown in Figure 4, retaining more semantic information during pretraining significantly enhances in-distribution regression performance, consistent with Cor. 4.1. Conversely, under distribution shift scenarios, accurately identifying invariant semantic variables is essential for robust out-of-distribution generalization. In this setting, introducing appropriate selection or perturbation biases effectively removes variables sensitive to distribution shift, supporting the result in Cor. 4.2. Additional results on the effects of perturbation bias are provided in § D.3.
@@ 5.2 MPI3d-Complex Dataset
The MPI3D-Complex dataset [18] consists real-world images captured in a controlled environment, with mutually independent, discrete latent factors: object color (color; 4 values), shape (shape; 4 values), size (size; 2 values), camera height (cam.; 3 values), background color (back.; 3 values), and object position along the horizontal (hori.; 40 values) and vertical (vert.; 40 values) axes. We designate the positional factors (hori., vert.) as image-specific, while the remaining attributes are treated as semantic variables. Text observations are generated based on the ground-truth semantic variables, with variation introduced via different selection and perturbation bias settings, and augmented with template-based text generation to simulate text-specific factors.
@@
Training is performed using the multimodal contrastive loss $\mathcal{L}_{\text{\tiny MMCL}}$ defined in Eq. 1, following the formulation in [12]. Performance is evaluated using the average Matthews Correlation Coefficient (MCC) across three random seeds, based on the prediction of all image latent variables from the learned image and text representations using linear and nonlinear MLP decoders, respectively. Further implementation details—including dataset statistics, bias configurations, text generation procedures, encoder architectures, and training hyperparameters—are provided in § E.1.
@@ Results.
Figure 5 presents the MCC results using nonlinear classifiers with learned representations. The findings demonstrate that, even with discrete latent variables, misaligned semantic variables across modalities, whether due to selection or perturbation biases, are systematically excluded from the representations (MCC $=0$). In contrast, unbiased semantics are well recovered, with MCC scores predominantly approaching 1 and all values $\geq 0.8$, reinforcing our theoretical findings. Further investigations into MCC using linear classifiers and ablation studies on encoder dimensionality are provided in § E.2.
@@ 5.3 Causal3DIdent Dataset
We conduct our experiments using the Causal3DIdent dataset, a semi-synthetic benchmark widely used in the causal representation learning literature [80, 66, 12, 76]. This dataset enables explicit enforcement of causal structures over latent variables. The generative process for images is governed by a set of 10 latent variables, comprising 3 discrete factors—object shape (shape), and object positions along the horizontal (x_pos) and vertical (y_pos) axes—and 7 continuous factors: object color (color); spotlight position (s_pos) and color (s_color); background color (b_color); and object rotation angles (alpha, beta, gamma). We treat the rotation angles (alpha, beta, gamma) as image-specific variables, and the remaining factors as semantic variables, which follow a predefined causal structure illustrated in Figure 14.
@@
We synthesize 80,000 samples for MMCL pretraining, 10,000 samples for training classifiers or regressors, and additional 10,000 samples for evaluating predictive performance. Evaluation metrics include $R^{2}$ for continuous latent variables and MCC for discrete latent variables. Further details on dataset specifications, encoder architectures, and training parameters, are provided in § F.1.

### v2

@@ 3.2 Model Assumptions for Theoretical Analysis
We now present the assumptions underlying our theoretical analysis, based on the proposed LVM:
@@
Remark 3.1.
Note that Eq. 5 essentially implies that, in each $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$, only a subset of semantic variables $A$ undergoes perturbations, regardless of the underlying causal structure among latent semantic variables. The rationale is that latent semantic variables can only be modified indirectly by altering observations, rather than through direct intervention, unless the latent causal structure is fully identified. In the text modality, it occurs through the misassignment of certain content words to specific image semantics during the captioning process. Unlike a direct intervention, this misalignment does not propagate to descendant semantic variables, as illustrated in Figure 2.
4 Identifiability Results for the Proposed Latent Variable Model
In this section, we theoretically analyze how misalignment impacts the identifiability of latent semantic variables in the proposed LVM, within the MMCL framework (§ 4.1). Based on these results, we further provide practical insights into how misalignment should be addressed in real-world applications (§ 4.2). Detailed proofs of the theoretical results are provided in App. B. To begin, we restate the definition of block-identifiability (Von Kügelgen et al., 2021) in the context of our problem setting:
@@ 4 Identifiability Results for the Proposed Latent Variable Model
4.1 Identifiability of Latent Semantic Variables under Misalignment
@@
Let $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$ be image-text pairs drawn from the data-generating process described in § 3, where $\mathbf{x}$ is generated according to Eq. 3 and $\mathbf{t}_{\theta}$ is generated by Eq. 4. Suppose that Asms. 3.1 and 3.2 hold. Denote by $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ the subset of semantic variables that preserved without bias in the text, and define its dimension as $n=|\mathbb{I}_{\theta}|-|\mathbb{I}_{\rho}|$. Let $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n}$ be sufficiently flexible, smooth functions. Then, Minimizing the loss $\mathcal{L}_{\text{SymAlignMaxEnt}}$ in Eq. 2 over samples $(\mathbf{x},\mathbf{t}_{\theta})$ guarantees that $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ block-identify the semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ in the sense of Defn. 4.1.
Remark 4.1.
Thm. 4.1 formally establishes that, in the presence of misalignment, the unbiased semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ that are shared across modalities can be effectively recovered, up to a block-wise indeterminacy, by minimizing the MMCL objective. In contrast, components that are misaligned, specifically $\mathbf{s}_{\mathbb{I}_{\rho}}$ and $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$, are entirely excluded from the learned representations. We emphasize that this result holds regardless of any underlying latent graph structure among semantic variables. At its core, this result highlights the model’s capacity to focus exclusively on the aligned semantic aspects of the data. Furthermore, modality-specific, non-semantic factors, i.e., $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, are consistently discarded throughout the learning process. This further underscores the model’s ability to extract meaningful, cross-modal semantic information while filtering out semantically uninformative or noise-induced components. A detailed proof of Thm. 4.1 is provided in § B.2.
@@ 4.2 Insights into Misalignment for Practice
Let the selection bias be $\theta=2^{n_{s}}-1$ and the perturbation bias be $\rho=1$, such that the full set of semantic variables $\mathbb{I}_{\mathbf{s}}$ is selected, and the perturbable semantic subset is trivial, i.e., $\mathbb{I}_{\rho}=\emptyset$. Then, all semantic variables $\mathbf{s}$ are block-identified via smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n_{s}}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n_{s}}$, when minimizing $\mathcal{L}_{\text{SymAlignMaxEnt}}$.
Insight 4.1 (MMCL Pretraining on Large-Scale Data). Large-scale multimodal datasets (e.g., COCO (Lin et al., 2014), Conceptual Captions (Sharma et al., 2018), LAION-5B (Schuhmann et al., 2022)) often exhibit varying caption quality. Our analysis indicates that omitted or perturbed semantics are irretrievably lost in the learned representations, regardless of dataset size, although scale may mitigate sporadic misalignment by averaging its effects. Preserving a breadth of relevant semantic details is therefore crucial when pretraining foundation models, whose primary goal is to support diverse downstream tasks. As noted in Cor. 4.1, achieving this requires detailed and consistent annotation of image semantics. Consequently, improved caption control (Li et al., 2022; Fang et al., 2015; Ding et al., 2023) is essential to avoid blind spots in semantic coverage.
@@
Consider an OOD setting in which a subset of semantic variables, $\mathbb{I}_{inv}\subset\mathbb{I}_{\mathbf{s}}$, remains invariant between training and testing environments, while the remaining semantic variables, $\mathbb{I}_{var}=\mathbb{I}_{\mathbf{s}}\setminus\mathbb{I}_{inv}$, undergo distribution shifts. If the union of omitted and perturbable semantic variables under selection bias $\theta$ and perturbation bias $\rho$ coincides with the environment-sensitive subset, i.e., $\mathbb{I}_{var}=\mathbb{I}_{\theta}^{c}\cup\mathbb{I}_{\rho},$ then the invariant semantic variables $\mathbf{s}_{\mathbb{I}_{inv}}$ are block-identified via smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{|\mathbb{I}_{inv}|}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{|\mathbb{I}_{inv}|},$ by minimizing $\mathcal{L}_{\text{SymAlignMaxEnt}}$.
Insight 4.2 (Invariant Representation Learning). In tasks requiring robust OOD performance (e.g., domain generalization) (Rojas-Carulla et al., 2018; Arjovsky et al., 2019), semantic variables that are vulnerable to distribution shifts can undermine generalization. As noted in Cor. 4.2, misalignment may, counterintuitively, enhance robustness by selectively omitting or perturbing these vulnerable variables. This suggests that MMCL may offer a novel perspective on invariant representation learning (Peters et al., 2016; Krueger et al., 2021; Eastwood et al., 2022), as auditing and curating text is more precise and interpretable, since language is distilled from human knowledge.
@@ 5.1 Numerical Simulation
We synthesize numerical data following the generative process described in § 3. Specifically, we sample modality-specific variables $\mathbf{m}_{x}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{x}})$ and $\mathbf{m}_{t}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{t}})$, each of dimension 5, along with full semantic variables $\mathbf{s}\sim\mathcal{N}(0,\Sigma_{\mathbf{s}})$, of dimension 10. Potential causal dependencies among these variables are encoded in their covariance matrices $\Sigma_{(\cdot)}$. To simulate various misalignment scenarios, we progressively increase the strength of selection and perturbation biases. For selection bias, we incrementally define selected subsets $\mathbb{I}_{\theta}$ as $\{1\},\dots,[10]$. [^4] For any positive integer $n>1$, we use the notation $[n]$ to denote the set $\{1,2,\dots,n\}$. For perturbation bias, we similarly use increasing subsets of $\mathbb{I}_{\rho}$, ranging from the empty set to $[9]$, applying additive Gaussian noise with a probability 0.75 for each semantic dimension $i\in\mathbb{I}_{\rho}$. These biases are applied separately: when analyzing selection bias, we set $\mathbb{I}_{\rho}=\emptyset$; conversely, when analyzing perturbation bias, we fix $\mathbb{I}_{\theta}=[9]$. Generation functions $\mathbf{g}_{x}$ and $\mathbf{g}_{t_{\theta}}$ are instantiated as randomly initialized invertible MLPs. Two modality-specific MLP encoders are trained for 100,000 steps using the $\mathcal{L}_{\text{SymAlignMaxEnt}}$ loss defined in Eq. 2, setting the representation dimension equal to the unbiased semantic dimensions. Further details on the experimental setup are provided in § C.1.
We conduct two main experiments. (i) Identification of semantics: We predict each dimension of the true semantic variables from the learned representations using a lightweight MLP, reporting the predictive $R^{2}$ score on holdout data. Results are averaged over three random seeds for each setting. (ii) Downstream performance: We evaluate the pretrained representations obtained under various bias conditions on downstream tasks. Specifically, we construct four regression task labels $y_{1},y_{2},y_{3}$, and $y_{4}$ by applying complex nonlinear functions to subsets of the semantic variables: $[3]$, $[5]$, $[7]$, and $[9]$, respectively. For classification, we binarize $y_{2}$ to produce binary labels. To evaluate the OOD generalization performance of the trained classifiers, we introduce a distribution shift in the observations $\mathbf{x}$ by applying a heavy-tailed transformation to dimensions $\{9,10\}$. Further details on the task design are provided in § C.1.
@@
The results in Figure 3 show that, under the independent latent variable scenario, unbiased semantic variables are clearly block-identified ($R^{2}\approx 1$), whereas misaligned semantics due to selection bias are effectively discarded ($R^{2}\approx 0$). In the dependent latent variable scenario, some misaligned semantics become partially predictable, reflecting inherent mutual predictability among strongly dependent variables (Von Kügelgen et al., 2021; Yao et al., 2023). Modality-specific variables are consistently omitted from the representations across all settings. Similar effects are observed under perturbation bias, as illustrated in Figure 7. Notably, although our theoretical results hold only up to invertible mappings, simple linear regression already achieves high $R^{2}$ scores, as demonstrated in Figure 8. These findings consolidate the identifiability results in Thm. 4.1. Additional analyses on misassigned encoding dimensions and combined bias effects, are provided in § C.2.
@@
Downstream Performance.
As shown in Figure 4, retaining more semantic information during pretraining significantly enhances in-distribution regression performance, consistent with Cor. 4.1. Conversely, under distribution shift scenarios, accurately identifying invariant semantic variables is essential for robust out-of-distribution generalization. In this setting, introducing appropriate selection or perturbation biases effectively removes variables sensitive to distribution shift, supporting the result in Cor. 4.2. Additional results on the effects of perturbation bias are provided in § C.3.
@@ 5.2 MPI3d-Complex Dataset
Training is performed using the multimodal contrastive loss $\mathcal{L}_{\text{MMCL}}$ defined in Eq. 1, following the formulation in (Daunhawer et al., 2022). Performance is evaluated using the average Matthews Correlation Coefficient (MCC) across three random seeds, based on the prediction of all image latent variables from the learned image and text representations using linear and nonlinear MLP decoders, respectively. Further implementation details—including dataset statistics, bias configurations, text generation procedures, encoder architectures, and training hyperparameters—are provided in § D.1.
@@ Results.
Figure 5 presents the nonlinear MCC results with learned representations. The findings demonstrate that, even with discrete latent variables, misaligned semantic variables across modalities, whether due to selection or perturbation biases, are systematically excluded from the representations (MCC $=0$). In contrast, unbiased semantics are well recovered, with MCC scores predominantly approaching 1 and all values $\geq 0.8$, reinforcing our theoretical findings. Further investigations into MCC using linear classifiers and ablation studies on encoder dimensionality are provided in § D.2.
@@ 5.3 Causal3DIdent Dataset
We conduct our experiments using the Causal3DIdent dataset, a semi-synthetic benchmark widely used in the causal representation learning literature (Zimmermann et al., 2021; Von Kügelgen et al., 2021; Daunhawer et al., 2022; Yao et al., 2023). This dataset enables explicit enforcement of causal structures over latent variables. The generative process for images is governed by a set of 10 latent variables, comprising 3 discrete factors—object shape (shape), and object positions along the horizontal (x_pos) and vertical (y_pos) axes—and 7 continuous factors: object color (color); spotlight position (s_pos) and color (s_color); background color (b_color); and object rotation angles (alpha, beta, gamma). We treat the rotation angles (alpha, beta, gamma) as image-specific variables, and the remaining factors as semantic variables, which follow a predefined causal structure illustrated in Figure 14.
@@
Evaluation metrics include $R^{2}$ for continuous latent variables and MCC for discrete latent variables. Further details on experimental setup, including dataset specifications, encoder architectures, and training parameters, are provided in § E.1.